Dunstreet's Department Store would like to develop an inventory ordering policy of a 95 percent probability of not stocking out. To illustrate your recommended procedure, use as an example the ordering policy for white percale sheets. Demand for white percale sheets is 3,400 per year. The store is open 365 days per year. Every two weeks (14 days) inventory is counted and a new order is placed. It takes 9 days for the sheets to be delivered. Standard deviation of demand for the sheets is seven per day.

There are currently 120 sheets on hand. How many sheets should you order? (Use Excel's NORMSINV() function to find the correct critical value for the given α-level. Do not round intermediate calculations. Round "z" value to 2 decimal places and final answer to the nearest whole number.)
Answer #1

Order interval(T) = 14 days

Lead time (L) = 9 days

Annual demand (D) = 3400 sheets

Number of days per year = 365 days

Average daily demand(d) = D/number of days per year = 3400/365 = 9.315 sheets

Standard deviation of daily demand(\sigma d) = 7 sheets

Service level = 95%

With a service level of 95% the value of Z = 1.65

On hand inventory(I) = 120 sheets

So,order quantity = [d(T+L)] + [Z X \sigmad X √(T+L)] - I

= [9.315(14+9)] + [1.65 X 7 X √(14+9)] - 120

= (9.315 X 23) + (1.65 X 7 X √23) - 120

= 214.245 + (1.65 x 7 x 4.80) - 120

= 214.245 + 55.44 - 120

= 269.685 - 120

= 149.685 or rounded to 150 sheets

So 150 sheets should be ordered